OpenStudy (anonymous):

the product of a and b is equal to x, and the sum of a and b is equal to y. If x and y are positive numbers, what is 1\[\div a + 1\div b\] in terms of x and y

jhonyy9 (jhonyy9):

a*b=x a+b=y - than 1/a + 1/b = (a+b)/ab = y/x